# Colorful Minors

### v1

@@ 1.1 Our results
Let and let be a -colorful graph problem that is both folio-representable and definable by some CMSO-formula Then, there exists an algorithm that, given a -colorful graph decides whether in time
Given folio-representability as a property of our problems, in theorem 1.9 plays the same role as treewidth does in Courcelle’s Theorem. Moreover, by theorem 1.6, corresponds to the largest integer such that contains a -segregated grid as a colorful minor. By the pigeon hole principle this means that is equivalent to the largest such that contains a -segregated grid for some color as a colorful minor.
For our second result, we consider a fragment of denoted as In this logic, quantification is restricted to sets of bounded bidimensionality, and a special predicate is available, asserting the existence of vertex-disjoint paths connecting the pairs It was shown in [81] that if a graph class is definable in then membership in can be decided in time where is the Hadwiger number of For more on the expressibility of this logic, see [81, 30, 36, 37].
We define the colorful Hadwiger number of a -colorful graph as the maximum for which contains the -colorful rainbow as a colorful minor and we denote it by As before, we define the strict colorful Hadwiger number of as
We prove the following.
@@
Let and let be a -colorful graph problem that is folio-representable and definable by some -formula Then, there exists an algorithm that, given a -colorful graph decides whether in time
@@
We emphasize that the applicability of Theorems 1.9 and 1.10, based on the structural theorems Theorems 1.5 and 1.6 respectively, exceeds the scope of Courcelle’s Theorem and the results of [81]. There are -colorful graphs for which (resp. ) is small, while the treewidth (resp. Hadwiger number) of is unbounded. We derive Theorems 1.9 and 1.10 in section 8 as special cases of a more general result applicable to optimization problems (theorem 8.4). In fact, in section 8, we see these theorems as corollaries of broader statements (theorem 8.9 and theorem 8.10) formulated in terms of optimization problems where we additionally display the parameter dependencies hidden in the -notation.
In appendix A, we present several problems on colorful graphs to which our framework applies – and for which no comparable algorithmic results were previously known.
1.2 Related work
@@
The most extensively studied problem on -colorful graphs is the Steiner Tree problem, which was shown to be solvable in polynomial time when the restrictive treewidth is bounded [45], and when a rainbow is excluded [41]. To the best of our knowledge, no combinatorial or algorithmic study has been undertaken on -colorful graphs as a general and unifying combinatorial framework.
1.3 Organization of the paper
section 2 introduces all preliminary concepts we require throughout the paper. section 2.3 particularly introduces key concepts from from graph minor theory for handling the structure of -minor-free graphs. Moreover, section 2.4 establishes our well-quasi-ordering result. Afterwards, section 3, section 4, and section 5 concern our structural results. section 3 deals with the proof of the full version of theorem 1.5, namely theorem 3.6, on the structure of colorful graphs excluding a rainbow clique as a colorful minor. In section 3.3, we prove a technical lemma we dub the “clique-compression lemma” (lemma 3.9) which will be crucial for the arguments necessary to establish the algorithm for colorful minor checking in section 7. section 4 proceeds with the full version of theorem 1.7, namely theorem 4.3, on the structure of colorful graphs excluding a rainbow grid as a colorful minor. The last section on structure results, section 5, deals with the structure of graphs excluding segregated grids, in the form of theorem 5.3, which is the full version of theorem 1.6. section 6 concerns our results on the Erdős-Pósa property for colorful graphs. section 6.3 deals with the connected case of the positive part of theorem 1.8, while in section 6.4 we lift this result to the disconnected case. Finally in section 6.5, we complete this section with the negative part of theorem 1.8. Going into section 8, we discuss algorithmic applications of our colorful minor framework and we present our two AMTs, namely theorem 8.9 and theorem 8.10, the full versions of theorem 1.9 and theorem 1.10 respectively. In section 9 we discuss possible directions for future research. Last but not least, in Appendix A we present a (non-exhaustive) list of problems in which our AMTs apply. Appendix B concerns the proof of a technical lemma we deemed best to move to the Appendix.
@@ 2.1 Minors
2.2 Highly linked sets and tangles.
@@ 2.3 Graph minor structure
We continue by briefly introducing some key-concepts for handling the general structure of -minor-free graphs. The definitions we introduce here act as strict simplification of the notions and terminology used in most proofs of the Graph Minor Structure Theorem by Robertson and Seymour. For more in-depth discussions on this topic, the reader is referred to [72, 52, 38]. For the purposes of this paper, the simplified definitions are fully sufficient as we are mostly using the corresponding theorems as black boxes. Most of the notions below are adapted from [51].
@@
If is a path, we say that is grounded if has at least one edge. In case is a cycle we say that is grounded if is a simple777That is, if has at least three vertices and no parallel edges. cycle. If is grounded, we call the subset of that is the union of together with the union of the drawings of all edges of the trace of Notice that the trace of is a simple curve in case is a path and a closed curve in case is a cycle.
@@ 2.4 Well-Quasi Ordering
This is all we need to deduce theorem 1.1.
Proof of Theorem 1.1..
@@
Then there must exist such that and satisfy the outcome of proposition 2.6 witnessed by the map Now observe that also witnesses that is a minor of and, in particular, for all Hence, is a colorful minor of which contradicts the assumption that is an anti-chain. ∎
@@
Let be a non-negative integer and let be a class of -colorful graphs that is closed under taking colorful minors. We define the (colorful minor) obstruction set of denoted by as the set of all colorful minor minimal -colorful graphs that are not members of Notice that the obstruction set of is an anti-chain for the colorful minor relation, therefore, from theorem 1.1 we have the following.
@@ 3 Excluding a rainbow clique
In this section we discuss the structure of colorful graphs excluding a fixed rainbow clique as a colorful minor. For this, first we establish a general tool that allows us to connect several sets of vertices to a fixed target set with many disjoint paths, or to find a small separator that separates at least one of the from From here, we make use of several useful results of Robertson and Seymour from [79] to establish the following key theorem of this section:
@@
With theorem 3.1, we are ready to prove two core follow-up results. The first of those results is the structure theorem for colorful graphs without rainbow clique minors, i.e. theorem 1.5. The second result is a further strengthening of theorem 3.1 we call the “clique-compression lemma” (see lemma 3.9) which will be useful for establishing an irrelevant vertex rule to be used in the algorithmic results in section 7. Informally speaking, the clique-compression lemma says that, in the presence of a large clique-minor in a colorful graph with colors, one can partition the set into two sets and and find a small set such that the -big component of only has vertices with colors from and contains a large “partially rainbow” clique minor in the sense that every branchset of contains all colors from with
@@ 3.1 Finding a rainbow clique
Within our proofs, in this section and beyond, we will repeatedly make use of a variant of proposition 2.1 that allows us to link a family of a bounded number of vertex sets equally into a single set or separate at least one of the from by deleting a small set of vertices.
@@
For all positive integers every graph and every collection of vertex sets one of the following holds:
@@
Let and be given. We construct an auxiliary graph as follows. For every we introduce a set of new vertices and for each we make adjacent to all vertices of We then apply proposition 2.1 to the graph with sets and to either find a --linkage of order or a --separator of order at most in
Let us first assume the case where proposition 2.1 returns the linkage Then every vertex of must be an endpoint of some path in For each let the collection of paths obtained from a path with one endpoint in by deleting this endpoints from It follows that and is an --linkage. Hence, this case produces a linkage as desired.
So now assume proposition 2.1 returns a --separator of order at most in Let be the set of all such that there is no --path in All we have to do is to show that is non-empty. Towards this goal let be the collection of all such that is non-empty. Since and we know that Now observe that there cannot be a such that there is an --path in To see this assume that such a path, let us call it exists for some Then this path must also exist in However, since there is some and has one endpoint in the path is a --path in This contradicts the assumption that is a --separator in Thus, the set is the set required by the second outcome of our assertion and the proof is complete. ∎
@@
Notice that proposition 3.3 can be realized in polynomial time, allowing for an algorithmic version. From here on, all of our proofs are constructive and thus, all of them admit polynomial-time algorithms that yield the respective outcomes.
@@
From here on we may assume that there does not exist a separation in such that and there is some such that In this situation we may apply proposition 3.3 to and the graphs It follows that there exists some together with connected and pairwise vertex-disjoint subgraphs of such that
@@
Proof of Theorem 3.1..
Let We may now apply lemma 3.5.
Suppose lemma 3.5 returns a separation of of order at most together with a non-empty set such that and there is some for which It follows that is restricted. Moreover, the unique -big component of must be fully contained in Hence, with we have found the desired set.
Hence, we may assume that lemma 3.5 returns pairwise vertex-disjoint subgraphs where and
@@ 3.2 Colorful graphs without large rainbow clique minors
We now have everything in place to prove theorem 1.5. In fact, we will prove a slightly stronger statement which will then imply theorem 1.5.
@@
Our proof of theorem 3.6 proves a slightly stronger variant that allows for a straightforward induction.
@@
We begin by declaring the functions involved. Let be the constant from proposition 2.3.
Notice that the required bounds on all four functions follow directly from proposition 2.5 and the above definitions.
@@
We now call proposition 2.2 for and which has two possible outcomes. Either, is -linked, or there is a -balanced separator of size at most for We treat these two cases separately.
@@
Since we may now apply proposition 2.5 to and obtain a -wall such that This allows us to apply proposition 2.5 to asking for a clique minor of size There are two possible outcomes:
@@
In this case, we apply theorem 3.1 to in Since, by our assumption, does not contain a rainbow as a colorful minor, the second outcome of theorem 3.1 must hold. That is, we obtain a set of size at most such that the -big component of is restricted.
@@
3.3 The clique-compression lemma
The main goal of this subsection is to provide a further refinement of theorem 3.1. While theorem 3.1 already provides a powerful way to understand clique minors in graphs without rainbow cliques, we need a much stronger version to realize our algorithmic aspirations in section 7.1. That is, we want to show that, in the presence of a very large clique minor, we can always find a subset of colors together with a bounded size set such that (i) there is a big clique minor where each branchset meets every color in and this clique minor is separated by from all vertices whose palette holds a color that does not belong to
Corollary 3.8.
Let and be non-negative integers with Let be a -colorful graph, be a set of size and be a minor model of in Then one of the following is true.
i)
There exists a colorful minor model of a rainbow in such that for all and is a truncation of or
ii)
Proof.
The claim follows immediately from applying theorem 3.1 to ∎
Lemma 3.9.
Let and be non-negative integers with Let be a -colorful graph and be a minor model of in
Then one of the following is true:
i)
ii)
there exists a near888That is, and but we allow one of the to be empty. partition of with such that the following statements hold:
a)
There is a set of size at most such that, if is the -big component of then
b)
there exists a minor model of in such that is a truncation of where and
c)
if is the restriction of to and then is a colorful minor model of a rainbow in
a)
Proof.
We define the following recursive function for all positive integers :
Notice first that
We begin by applying theorem 3.1 to and The first outcome of theorem 3.1 is the same as the first outcome of the assertion, so in this case we would be done. Hence, we may assume that the second outcome of theorem 3.1 holds. This provides us with a set of size at most such that the -big component of is restricted. Let be the minor model of a complete graph obtained from by removing all with Moreover, let be the -big component of and let Then is a minor model of a complete graph in Finally, let be the restriction of to and Then is a colorful graph.
We prove the following claim by induction on
Claim 1. Whenever we are given a graph a set of size at most and a -colorful graph containing a minor model of where we can find one of the following outcomes.
i)
There exists a colorful minor model of a rainbow in such that is a truncation of where or
ii)
Proof of Claim 1. Claim 1 follows almost immediately from corollary 3.8. To see this notice that
Hence, corollary 3.8 either provides a rainbow clique of size or a set of size at most such that the -big component of is restricted as required.
We start with and apply Claim 1 for and as defined above. If Claim 1 returns, for some a set of size at most via its second outcome, we may define Moreover, if is the -big component of then we know that is also the -big component of In this case we set and define to be the minor model of a complete graph obtained from by removing all that have a non-empty intersection with To be able to repeatedly apply Claim 1, it suffices to show that for all As before, this easily follows by induction on
Finally, notice that in the case where we have that In this case we may set and since the -big component of cannot have a single color. This implies that, with
that is indeed a colorful minor model of a rainbow in of the size required by the second outcome of our assertion. Hence, after at most applications of Claim 1, we find the required rainbow minor model together with the required set and our proof is complete. ∎
4 Excluding a rainbow-grid
We now move on from rainbow cliques to rainbow grids: In this section, we explain how theorem 1.7 can be deduced from a powerful variant of proposition 2.5 by Paul, Protopapas, Thilikos and Wiederrecht [69, 68].

### v2

@@ 1.1 Our results
Let and let be a -colorful graph problem that is both folio-representable and definable by some CMSO-formula Then, there exists an algorithm that, given a -colorful graph decides whether in time where is the maximum monodimensionality of a color in .
From theorem˜1.3, the maximum monodimensionality of a color in a -colorful graph is equivalent to the maximum torso treewidth of over all colors Notice that theorem˜1.8 can be restated by setting to be this equivalent quantity, which plays the same role for colorful graphs as treewidth does for graphs in Courcelle’s Theorem.
For our second result, we consider a fragment of denoted as In this logic, quantification is restricted to sets of bounded bidimensionality, and a special predicate is available, asserting the existence of vertex-disjoint paths connecting the pairs It was shown by Sau, Stamoulis, and Thilikos [81] that if a graph class is definable in then membership in can be decided in time .
The Hadwiger number of a color in a -colorful graph is the maximum for which contains the rainbow -clique as a colorful minor. We prove the following.
@@
Let and let be a -colorful graph problem that is folio-representable and definable by some -formula Then, there exists an algorithm that, given a -colorful graph decides whether in time , where is the maximum Hadwiger number of a color of .
We emphasize that the applicability of Theorems˜1.8 and 1.9, based on the structural theorems Theorems˜1.2 and 1.3 respectively, exceeds the scope of Courcelle’s Theorem and the results of [81]. The generality of our extension is displayed by the fact that, for every , there are classes of -colorful graphs for which the maximum monodimensionality/Hadwiger number of a color in is bounded while the treewidth/Hadwiger number of is unbounded.
This is a strong indication that, for problems on colorful graphs, the appropriate structural parameterizations toward deriving (meta-)algorithmic results should take into account the way the colors are distributed in the inputs.
@@
We derive Theorems˜1.8 and 1.9 in section˜7 as special cases of a more general result applicable to optimization problems (theorem˜7.4). In fact, in section˜7, we see these theorems as corollaries of broader statements (theorem˜7.9 and theorem˜7.10) formulated in terms of optimization problems where we additionally display the parameter dependencies hidden in the -notation.
1.2 Brief insight into our approach
The general approach to proving Robertson-Seymour-style structure theorems typically constists of establishing a local structure theorem with respect to a wall and then applying a well-known technique, originating from the work of Robertson and Seymour [78], that turns the local statement into a global one based on tree-decompositions. Our structure theorems 1.2, 1.3, and 1.4 follow this general paradigm.
In the uncolored setting, proving a local structure theorem with respect to a wall usually proceeds as follows: Starting from a large wall in the graph , one either finds
a.i)
a large clique-minor that is highly connected to or
a.ii)
a structural decomposition of relative to typically of a topological nature, such as a flat wall or, more generally, a near embedding of in some surface with additional properties tailored to the theorem at hand.
For our results, handling outcome (a.ii) relies on sophisticated topological arguments stemming from the Graph Minor Structure Theorem [72] and its subsequent improvements [53, 38] as well as more recent refinements [83, 68, 69]. While technically demanding, this part largely builds on and adapts existing machinery to deal with annotation in near embeddings [69], except for theorem˜1.3 which uses different techniques which we comment on in the next paragraph.
Rainbow clique-minors and a multicolored Menger-type theorem.
The genuinely new difficulties and contributions arise in outcome (a.i). In the colored settings, the mere presence of a large clique-minor provides little information on the structure of the colorful graph with respect to the colored vertices. For our purposes, it is therefore essential to understand how well the colored vertices are connected to this clique-minor. Concretely, given a -colorful graph containing a large clique-minor, we aim to establish the following dichotomy:
b.i)
there exists a small set of vertices and a color such that the component of containing most of the clique-minor is free of vertices of color or
b.ii)
contains a large rainbow clique-minor that is highly connected to the original clique-minor.
This results, formalized in theorem˜3.1, is proved in section˜3.1 and constitutes a central new ingredient underpinning all our structural theorems.
In the context of the Disjoint Paths problem, a similar result was established by Robertson and Seymour [79] for -colorful graphs with a bounded number of colored vertices. Our setting is substantially more general: we allow -colorful graphs with an arbitrary number of colored vertices. Our proof involves a one-to-many strengthening of Menger’s Theorem, stated as lemma˜3.2, which may be of independent interest. For all positive integers , given vertex sets and in a graph either
c.i)
there exist disjoint paths in with exactly paths linking to or
c.ii)
there exists an index and a vertex set of size less than that separates from
Interestingly, our results on the Erdős-Pósa property for colorful minors imply that even a two-to-two strengthening of Menger’s theorem cannot hold in general, not even approximately (see section˜8.1). Consequently, our structural results are in a sense “tight”, as we cannot expect to replace condition (b.i) with a stronger alternative.
It is noteworthy that in the local version, theorem˜5.2, of theorem˜1.3, outcome (a.ii) also relies on the one-to-many strengthening of Menger above to control how well-connected the colored vertices are to a given flat wall in the graph.
Annotations and the Erdős-Pósa property.
Erdős-Pósa-type dualities in graph minors and beyond are intrinsically related to structural characterization of restricted classes of graphs, often in terms of “global” decomposition theorems based on tree-decompositions. A prime example of this interaction is the celebrated characterization of the proper minor-closed classes which have the Erdős-Pósa property for minors, shown by Robertson and Seymour [76] to be precisely those of bounded treewidth.
Crucial to this duality is showing that graph classes of unbounded treewidth necessarily fail to satisfy the Erdős–Pósa property. By the grid theorem, such classes are precisely those containing all planar graphs. Consequently, this reduces to understanding why the two Kuratowski minors and themselves do not admit the Erdős–Pósa property. In [76], Robertson and Seymour address this by constructing families of graphs based on and that explicitly violate the Erdős–Pósa property. These constructions are inherently topological: The resulting graphs embed minimally in the torus or the projective plane, but not in the sphere. At a high level, the argument exploits the additivity of Euler-genus under disjoint union [63], which implies that repeatedly combining non-planar graphs forces unbounded Euler-genus.
In the colorful setting, we initially follow the same high-level strategy. We identify torso treewidth as the correct parameter governing the Erdős–Pósa property for colorful minors. Utilizing our grid theorem for torso treewidth (theorem˜1.3), we identify the set of -colorful graphs and for the reverse direction of theorem˜1.5, we have to show that each -colorful graph in violates the Erdős–Pósa property for colorful minors.
At this point, however, the classical approach breaks down. Unlike the uncolored case, most graphs in are planar with particular assumptions on the topology induced by the colored vertices. As a consequence, the standard arguments are no longer applicable, as they fundamentally rely on non-planar embeddings.
Instead, our arguments crucially exploit the distribution of colored vertices within these graphs. We show that the topology induced by this distribution—rather than by the underlying graph alone—can still be leveraged to construct families that violate the Erdős–Pósa property. Once more, this highlights the importance of the arrangement of colored vertices within the underlying graph.
1.3 Related work and organization of the paper
Related work.
@@ 1.3 Related work and organization of the paper
The most extensively studied problem on -colorful graphs is the Steiner Tree problem, which was shown to be solvable in polynomial time when the torso treewidth is bounded [46], and when a rainbow is excluded [41]. To the best of our knowledge, no combinatorial or algorithmic study has been undertaken on -colorful graphs as a general and unifying combinatorial framework.
Organization of the paper.
section˜2 introduces all preliminary concepts we require throughout the paper. section˜2.3 particularly introduces key concepts from from graph minor theory for handling the structure of -minor-free graphs. Moreover, section˜2.4 establishes our well-quasi-ordering result and section˜2.5 the reduction for colorful minor checking. Afterwards, section˜3, section˜4, and section˜5 concern our structural results. section˜3 deals with the proof of the full version of theorem˜1.2, namely theorem˜3.6, on the structure of colorful graphs excluding a rainbow clique as a colorful minor. section˜4 proceeds with the full version of theorem˜1.4, namely theorem˜4.3, on the structure of colorful graphs excluding a rainbow grid as a colorful minor. The last section on structure results, section˜5, deals with the structure of graphs excluding segregated grids, in the form of theorem˜5.3, which is the full version of theorem˜1.3. section˜6 concerns our results on the Erdős-Pósa property for colorful graphs. section˜6.3 deals with the connected case of the positive part of theorem˜1.5, while in section˜6.4 we lift this result to the non-connected case. Finally in section˜6.5, we complete this section with the negative part of theorem˜1.5. Going into section˜7, we discuss algorithmic applications of our colorful minor framework and we present our two AMTs, namely theorem˜7.9 and theorem˜7.10, the full versions of theorem˜1.8 and theorem˜1.9 respectively. In section˜8 we discuss possible directions for future research. Last but not least, in Appendix A we present a (non-exhaustive) list of problems in which our AMTs apply.
@@ 2.1 Minors
2.2 Highly linked sets and tangles
@@ 2.3 Graph minor structure
We continue by briefly introducing some key concepts for handling the general structure of -minor-free graphs. The definitions we introduce here act as a strict simplification of the notions and terminology used in most proofs of the Graph Minor Structure Theorem by Robertson and Seymour. For more in-depth discussions on this topic, the reader is referred to [72, 53, 38]. For the purposes of this paper, the simplified definitions are fully sufficient as we are mostly using the corresponding theorems as black boxes. Most of the notions below are adapted from [52].
@@
If is a path, we say that is grounded if has at least one edge. In case is a cycle we say that is grounded if is a simple999That is, if has at least three vertices and no parallel edges. cycle. If is grounded, we call the subset of that is the union of together with the union of the drawings of all edges of the trace of Notice that the trace of is a simple curve in case is a path and a closed curve in case is a cycle.
@@ 2.4 Well-Quasi Ordering
This is all we need to deduce theorem˜1.1.
Proof of Theorem˜1.1..
@@
Then there must exist such that and satisfy the outcome of proposition˜2.6 witnessed by the map Now observe that also witnesses that is a minor of and, in particular, for all Hence, is a colorful minor of which contradicts the assumption that is an anti-chain. ∎
@@
Let be a non-negative integer and let be a class of -colorful graphs that is closed under taking colorful minors. We define the (colorful minor) obstruction set of denoted by as the set of all colorful minor minimal -colorful graphs that are not members of Notice that the obstruction set of is an anti-chain for the colorful minor relation, therefore, from theorem˜1.1 we have the following.
@@
2.5 Colorful minor checking
We now proceed with the proof of theorem˜1.6. Having established that the colorful minor relation is indeed a well-quasi-order, we next show that testing whether some -colorful graph is a colorful minor of a -colorful graph can be done efficiently. Towards proving theorem˜1.6 we require the fast minor testing algorithm of Korhonen, Pilipczuk, and Stamoulis [56]. Indeed, we need the rooted variant of their algorithm.
A tuple where for all is called a rooted graph and the are called the roots of . Notice that one may interpret a rooted graph with roots as an -colorful graph where for each . Indeed, to each rooted graph we associate a corresponding colorful graph such that for all . We say that a rooted graph is a rooted minor of a rooted graph if is a colorful minor of .
Proposition 2.8 (Korhonen, Pilipczuk, Stamoulis [56]).
There exists a function and an algorithm that takes as input two rooted graph and and decides if is a rooted minor of in time .
The proof of theorem˜1.6 is a simple reduction of the colorful minor checking problem to the rooted minor checking problem.
Let be a -colorful graph. We denote by the rooted graph obtained from by introducing for every and a new vertex adjacent precisely to .
Moreover, let be another -colorful graph. We denote by the rooted graph obtained from by introducing, for every and a new vertex adjacent to all vertices of .
Notice that for a fixed -colorful graph , the number of roots in and are the same. As a convention, we will assume the vertices of to come with some linear ordering such that in the sequence of roots of , the vertex has precisely the same position as the vertex in the sequence of roots of . Under this convention, we now obtain the following lemma.
Lemma 2.9.
Let and be -colorful graphs. Then is a colorful minor of if and only if is a rooted minor of .
Proof.
Assume first that there exists a colorful minor model of in . In this case, for each and there exists a vertex with . Hence, in , is adjacent to . It follows that is a minor model witnessing that is indeed a rooted minor of .
Now assume that there exists a minor model of in . In this case, for each , must first of all be disjoint from all roots of as each must be mapped by the minor model to the root of . Moreover, for each , there is a vertex of adjacent to . It follows from the definition of that . Hence, must be a colorful minor model of in . ∎
With lemma˜2.9 it now suffices to observe that both and can be computed from and in time linear in . Moreover, the number of roots of is at most and this is also an upper bound on the number of vertices has in excess of itself. Finally, the number of edges in is at most . So we may compute and run the algorithm from proposition˜2.8 to check for as a rooted minor without any additional impact on its running time in terms of and .
@@ 3 Excluding a rainbow clique
In this section we discuss the structure of colorful graphs excluding a fixed rainbow clique as a colorful minor. For this, first we establish a general tool that allows us to connect several sets of vertices to a fixed target set with many disjoint paths, or to find a small separator that separates at least one of the from From here, we make use of several useful results of Robertson and Seymour from [79] to establish the following key theorem of this section.
@@
Establishing theorem˜3.1 is the key towards our first structure theorem for colorful graphs excluding a rainbow clique minor, i.e. theorem˜1.2. Robertson and Seymour in [79], prove a theorem akin to theorem˜3.1 that however works in the setting of -colorful graphs with a bounded number of colored vertices. In the following section, we adapt this result to work for colors and an arbitrary number of colored vertices to suit our needs.
@@ 3.1 Finding a rainbow clique
Within our proofs, in this section and beyond, we will repeatedly make use of a variant of proposition˜2.1 that allows us to link a family of a bounded number of vertex sets equally into a single set or separate at least one of the from by deleting a small set of vertices.
@@
For all positive integers every graph and every collection of (not necessarily disjoint) vertex sets one of the following holds:
@@
Let and be given. We construct an auxiliary graph as follows. For every we introduce a set of new vertices and for each we make adjacent to all vertices of We then apply proposition˜2.1 to the graph with sets and to either find a --linkage of order or a --separator of order at most in
Let us first assume the case where proposition˜2.1 returns the linkage Then every vertex of must be an endpoint of some path in For each let the collection of paths obtained from a path with one endpoint in by deleting this endpoints from It follows that and is an --linkage. Hence, this case produces a linkage as desired.
So now assume proposition˜2.1 returns a --separator of order at most in Let be the set of all such that there is no --path in All we have to do is to show that is non-empty. Towards this goal let be the collection of all such that is non-empty. Since and we know that Now observe that there cannot be a such that there is an --path in To see this assume that such a path, let us call it exists for some Then this path must also exist in However, since there is some and has one endpoint in the path is a --path in This contradicts the assumption that is a --separator in Thus, the set is the set required by the second outcome of our assertion and the proof is complete. ∎
@@
Notice that proposition˜3.3 can be realized in polynomial time, allowing for an algorithmic version. From here on, all of our proofs are constructive and thus, all of them admit polynomial-time algorithms that yield the respective outcomes.
@@
From here on we may assume that there does not exist a separation in such that and there is some such that In this situation we may apply proposition˜3.3 to and the graphs It follows that there exists some together with connected and pairwise vertex-disjoint subgraphs of such that
@@
Proof of Theorem˜3.1..
Let We may now apply lemma˜3.5.
Suppose lemma˜3.5 returns a separation of of order at most together with a non-empty set such that and there is some for which It follows that is restricted. Moreover, the unique -big component of must be fully contained in Hence, with we have found the desired set.
Hence, we may assume that lemma˜3.5 returns pairwise vertex-disjoint subgraphs where and
@@ 3.2 Colorful graphs without large rainbow clique minors
We now have everything in place to prove theorem˜1.2. In fact, we will prove a slightly stronger statement which will then imply theorem˜1.2.
@@
Our proof of theorem˜3.6 proves a slightly stronger variant that allows for a straightforward induction.
@@
We begin by declaring the functions involved. Let be the constant from proposition˜2.3.
Notice that the required bounds on all four functions follow directly from proposition˜2.5 and the above definitions.
@@
We now call proposition˜2.2 for and which has two possible outcomes. Either, is -linked, or there is a -balanced separator of size at most for We treat these two cases separately.
@@
Since we may now apply proposition˜2.3 to and obtain a -wall such that This allows us to apply proposition˜2.5 to asking for a clique minor of size There are two possible outcomes:
@@
In this case, we apply theorem˜3.1 to in Since, by our assumption, does not contain a rainbow as a colorful minor, the second outcome of theorem˜3.1 must hold. That is, we obtain a set of size at most such that the -big component of is restricted.
@@
4 Excluding a rainbow grid
We now move on from rainbow cliques to rainbow grids: In this section, we explain how theorem˜1.4 can be deduced from a powerful variant of proposition˜2.5 by Paul, Protopapas, Thilikos and Wiederrecht [69, 68].